integration rootsquare cos2A/sinA=

Grade:12th pass

1 Answers

Nishant Vora IIT Patna
9 years ago

Hello student, Please find the answer to your question below

I = \int \frac{cos2A}{sinA}dA I = \int \frac{1-2sin^{2}A}{sinA}dA I = \int cosecAdA - 2\int sinAdA I = \int cosecAdA - 2(-cosA) I = \int \frac{1+tan^{2}\frac{A}{2}}{2tan\frac{A}{2}}dA + 2cosA I = \int \frac{sec^{2}\frac{A}{2}}{2tan\frac{A}{2}}dA + 2cosA I = log(tan\frac{A}{2})+ 2cosA+c